# Clarifying the effect of mean subtraction on Dynamic Mode Decomposition

### v5

@@ D.6 Delay embedding can generate a Koopman invariant dictionary
The preceding arguments, taken in light of the derivation in Section A.1 that gives Eq. 24, suggest that $\tilde{\mathbf{C}}_{\rm d-delayed}$ is the representation of $\boldsymbol{\mathbf{\psi}}_{\rm d-delayed}$ in the eigen-basis $\boldsymbol{\mathbf{\phi}}$:
@@
With this machinery in place, a classical argument from systems theory, reviewed here as Lemma D.28, shows that $r-1$ delays are sufficient for $\boldsymbol{\mathbf{\psi}}_{\rm d-delayed}$ to be Koopman invariant.
Lemma D.28 (Eigenvector test for Observability [8]).
@@
Proof D.29 (Proof of Proposition 4.10).
The assumption that $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct KEFs ensures, by Definition 3.8, that every column of $\tilde{\mathbf{C}}$ is non-zero. Hence, by Eq. 52, we find that every column of $\tilde{\mathbf{C}}_{\rm d-delayed}$ is also non-zero.
Consequently, we can apply Proposition A.4 for $\boldsymbol{\mathbf{\psi}}_{\rm d-delayed}$, which tells us that Koopman invariance of $\boldsymbol{\mathbf{\psi}}_{\rm d-delayed}$ can be established by showing that the columns of $\tilde{\mathbf{C}}_{\rm d-delayed}$ are linearly independent. This can, in turn, be proven by applying Lemma D.28 for the matrix pair $(\mathbf{\Lambda},\tilde{\mathbf{C}})$.
The eigenvectors of $\mathbf{\Lambda}$ are the standard basis vectors. The assumption (3.8) guarantees that no column of $\tilde{\mathbf{C}}$ is $\boldsymbol{\mathbf{0}}$. Hence, by Lemma D.28, $\mathrm{obs}(\mathbf{\Lambda},\tilde{\mathbf{C}})$ has full column rank. Now, $d\geq r-1$ ensures that the rows of $\mathrm{obs}(\mathbf{\Lambda},\tilde{\mathbf{C}})$ form a subset of the rows in $\tilde{\mathbf{C}}_{\rm d-delayed}$. Therefore, $\tilde{\mathbf{C}}_{\rm d-delayed}$ also has full column rank.
D.7 Relative distance to DFT
In $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ (Algorithm 3), DMD-DFT equivalence is defined in terms of a set equality (Definition 4.13). Since set equalities are not numerically well-defined, we develop a computationally tractable alternative dubbed Relative distance to DFT. Constructing this indicator necessitates unpacking Algorithm 3 and results in a characterization [^20] Analogous to Lemma D.1 that revolves around a vector equality.
Let $\boldsymbol{\mathbf{\mu}}_{\rm d-delayed}$ denote the temporal mean of $\mathbf{Z}_{\rm d-delayed}$:
$\boldsymbol{\mathbf{\mu}}_{\rm d-delayed}:=\mathbf{Z}_{\rm d-delayed}\frac{\boldsymbol{\mathbf{1}}_{\theta+1}}{\theta+1}.$
The mean-subtracted version of $\mathbf{Z}_{\rm d-delayed}$ would, then, read:
The resulting DMD model, $\boldsymbol{\mathbf{c}}^{*}[(\mathbf{Z}_{\rm d-delayed})_{\rm ms}]$, encodes the $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ eigen-values, $\{\,\check{\lambda}_{i}\,\}_{i=1}^{\theta}$. Since the latter is at the heart of Definition 4.13 being numerically ill-defined, we use the former to get the following reformulation:
Lemma D.30 (Recasting $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}\equiv\textrm{DFT}$ as a vector equality).
A necessary and sufficient condition for $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ to be equivalent [^21] As described in Definition 4.13. to a temporal DFT is that the $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ model, $\boldsymbol{\mathbf{c}}^{*}[(\mathbf{Z}_{\rm d-delayed})_{\rm ms}]$, coincide with $-\mathbf{1}_{\theta}$:
$d\boldsymbol{\mathbf{\mu}}\textrm{DMD}\equiv\textrm{DFT}~{}~{}\iff~{}~{}\boldsymbol{\mathbf{c}}^{*}[(\mathbf{Z}_{\rm d-delayed})_{\rm ms}]~{}=~{}-\boldsymbol{\mathbf{1}}_{\theta}.$
Proof D.31.
By Definition 4.13, the left-hand side can be expanded in terms of the $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ eigenvalues, $\{\,\check{\lambda}_{i}\,\}_{i=1}^{\theta}$.
Since $\{\,\check{\lambda}_{i}\,\}_{i=1}^{\theta}$ are defined in Algorithm 3 to be the eigenvalues of the Companion matrix $\mathbf{T}\left(\boldsymbol{\mathbf{c}}^{*}[(\mathbf{Z}_{\rm d-delayed})_{\rm ms}]\right)$, we have the desired conclusion.
Hence, the root-mean-squared error between $\boldsymbol{\mathbf{c}}^{*}[(\mathbf{Z}_{\rm d-delayed})_{\rm ms}]$ and $-\mathbf{1}_{\theta}$ can measure the deviation of $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ from DFT:
Definition D.32 (Relative distance to DFT).
(54) $\texttt{Relative distance to DFT}\,[\theta,d]:=\frac{\lVert\boldsymbol{\mathbf{c}}^{*}[(\mathbf{Z}_{\rm d-delayed})_{\rm ms}]+\mathbf{1}_{\theta}\rVert}{\sqrt{\theta}}$
Remark D.33.
By construction, Relative distance to DFT is non-negative. Yet, it can only have a maximum value of $1$. To see this, define $(\mathbf{X}_{\rm d-delayed})_{\rm ms}$ as the sub-matrix corresponding to the first $\theta$ columns of $(\mathbf{Z}_{\rm d-delayed})_{\rm ms}$:
Now, the proof of Theorem 4.1 tells us that,
$\boldsymbol{\mathbf{c}}^{*}[(\mathbf{Z}_{\rm d-delayed})_{\rm ms}]~{}=~{}\mathcal{P}_{\mathcal{R}\left(\,(\mathbf{X}_{\rm d-delayed})_{\rm ms}^{H}\,\right)}\left(-\boldsymbol{\mathbf{1}}_{\theta}\right).$
Substituting this in Definition D.32, and using the identity,
$I=\mathcal{P}_{\mathcal{N}\left(\,(\mathbf{X}_{\rm d-delayed})_{\rm ms}\,\right)}+\mathcal{P}_{\mathcal{R}\left(\,(\mathbf{X}_{\rm d-delayed})_{\rm ms}^{H}\,\right)},$
we have:
Relative distance to $\texttt{Relative distance to DFT}\,[\theta,d]{}$ $=~{}\frac{\lVert\boldsymbol{\mathbf{c}}^{*}[(\mathbf{Z}_{\rm d-delayed})_{\rm ms}]+\mathbf{1}_{\theta}\rVert}{\sqrt{\theta}}$
$=~{}\frac{\lVert\mathcal{P}_{\mathcal{R}\left(\,(\mathbf{X}_{\rm d-delayed})_{\rm ms}^{H}\,\right)}\left(-\boldsymbol{\mathbf{1}}_{\theta}\right)-\left(-\mathbf{1}_{\theta}\right)\rVert}{\sqrt{\theta}}$
@@ D.7 Relative distance to DFT
$=~{}\frac{\lVert\mathcal{P}_{\mathcal{N}\left(\,(\mathbf{X}_{\rm d-delayed})_{\rm ms}\,\right)}\left(-\boldsymbol{\mathbf{1}}_{\theta}\right)\rVert}{\lVert-\boldsymbol{\mathbf{1}}_{\theta}\rVert}$
$\leq~{}1.$
Hence, the term Relative distance to DFT is appropriate.

### v6

@@ E.6 Delay embedding can generate a Koopman invariant dictionary
The preceding arguments, taken in light of the derivation in Section B.1 that gives Eq. 36, suggest that $\tilde{\mathbf{C}}_{\rm d-delayed}$ is the representation of $\boldsymbol{\mathbf{\psi}}_{\rm d-delayed}$ in the eigen-basis $\boldsymbol{\mathbf{\phi}}$:
@@
With this machinery in place, a classical argument from systems theory, reviewed here as Lemma E.16, shows that $r-1$ delays are sufficient for $\boldsymbol{\mathbf{\psi}}_{\rm d-delayed}$ to be Koopman invariant.
Lemma E.16 (Eigenvector test for Observability [9]).
@@
Proof E.17 (Proof of Proposition 4.10).
According to Proposition B.3, Koopman invariance of $\boldsymbol{\mathbf{\psi}}_{\rm d-delayed}$ can be established by showing that the columns of $\tilde{\mathbf{C}}_{\rm d-delayed}$ are linearly independent. To begin with, $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct KEFs. So, Eq. 21 says every column of $\tilde{\mathbf{C}}$ is non-zero. Furthermore, Eq. 66 tells us that $\tilde{\mathbf{C}}$ comprises the first $m$ rows of $\tilde{\mathbf{C}}_{\rm d-delayed}$. Hence, every column of $\tilde{\mathbf{C}}_{\rm d-delayed}$ is also non-zero, which means $\boldsymbol{\mathbf{\psi}}_{\rm d-delayed}$ also lies in the non-redundant span of the same $r$ distinct KEFs. Consequently, we can apply Proposition B.3 to infer that Koopman invariance of $\boldsymbol{\mathbf{\psi}}_{\rm d-delayed}$ is equivalent to $\tilde{\mathbf{C}}_{\rm d-delayed}$ possessing full column rank.
Now, $d\geq r-1$ means the rows of $\tilde{\mathbf{C}}_{\rm d-delayed}$ are a superset of the rows of $\mathrm{obs}(\mathbf{\Lambda},\tilde{\mathbf{C}})$.
Additionally, Lemma E.16, applied to the matrix pair $(\mathbf{\Lambda},\tilde{\mathbf{C}})$, tells us that $\mathrm{obs}(\mathbf{\Lambda},\tilde{\mathbf{C}})$ has full column rank. To see this, note that the eigenvectors of $\mathbf{\Lambda}$ are the standard basis vectors. Furthermore, we have already seen that no column of $\tilde{\mathbf{C}}$ is $\boldsymbol{\mathbf{0}}$. In other words, no eigenvector of $\mathbf{\Lambda}$ lies in the null-space of $\tilde{\mathbf{C}}$. Hence, Lemma E.16 tells us that the columns of $\mathrm{obs}(\mathbf{\Lambda},\tilde{\mathbf{C}})$ are linearly independent.
Therefore, $\tilde{\mathbf{C}}_{\rm d-delayed}$ also has full column rank.
Appendix F Impact of low-rank regularization on the computation of Relative distance to DFT
Recall, from Remark 5.4, that the discrepancy between $(\mathbf{X}_{\rm d-delayed})_{\rm ms}$ and its low rank approximation $(\hat{\mathbf{X}}_{\rm d-delayed})_{\rm ms}$ determines whether we compute Relative distance to DFT  or just its upper bound. Specifically, Relative distance to DFT  can be computed if and only if said discrepancy is low. In practice, we let $\sigma_{\rm Tail}$ denote the norm of the discrepancy/approximation error i.e.,
@@ Appendix F Impact of low-rank regularization on the computation of Relative distance to DFT
$\sigma_{\rm Tail}~{}:=~{}\left\lVert(\mathbf{X}_{\rm d-delayed})_{\rm ms}-(\hat{\mathbf{X}}_{\rm d-delayed})_{\rm ms}\right\rVert,$
and deem the discrepancy to be low when $\sigma_{\rm Tail}\leq 10^{-15}$.
So, for each system considered in Section 5, we now scrutinize the variation of $\sigma_{\rm Tail}$ with model order $(\theta)$ and number of time delays $(d)$.
For all three LTI systems, Figure 4 tells us there is low discrepancy until we are well into the over-sampled regime ($\theta\geq 14$), thereby validating that Fig. 1 indeed depicts Relative distance to DFT  when $\theta<14$. Specifically, when we take only $3$ or $6$ time delays, $\sigma_{\rm Tail}$ is negligible throughout Fig. 4 and, hence, the concomitant trends in Fig. 1 are indeed those of Relative distance to DFT. However, when we excessively take $24$ time delays, $\sigma_{\rm Tail}$ becomes non-trivial from $\theta~{}=~{}14,16~{}\textrm{and}~{}21$ for $\textrm{LTI}_{1a},\textrm{LTI}_{1b}~{}\textrm{and}~{}\textrm{LTI}_{3}$ respectively. This departure from low discrepancy is conveniently reflected in Fig. 1 as secondary spikes. Therefore, the visualizations in Fig. 1 corresponding to $d=24$ depict Relative distance to DFT  only until said secondary spikes, beyond which they represent just an upper bound on Relative distance to DFT.
(a) $\textrm{LTI}_{1a}$
(b) $\textrm{LTI}_{1b}$
(c) $\textrm{LTI}_{3}$
Figure 4: Variation of $\sigma_{\rm Tail}$ corresponding to the studies visualized in Fig. 1. Given a choice of $(\theta,d)$, a low value of $\sigma_{\rm Tail}$ means we have accurately computed Relative distance to DFT  while a large value indicates that we have only calculated an upper bound. Hence, for all three systems, when we only take $3$ or $6$ time delays, Relative distance to DFT  is accurately computed for all model orders. In contrast, when $24$ time delays are taken, $\sigma_{\rm Tail}$ spikes at $\theta=14,16~{}\textrm{and}~{}21$ for $\textrm{LTI}_{1a},\textrm{LTI}_{1b}~{}\textrm{and}~{}\textrm{LTI}_{3}$ respectively, and remains high thereafter. Therefore, the concomitant computations in Fig. 1 for $d=24$ are only accurate until the aforementioned values of $\theta$. Beyond this, what is visualized in only an upper bound on Relative distance to DFT.
In the Van der Pol oscillator, we see that Fig. 5 displays a trend similar to those seen for the LTI systems in Fig. 4. The value of $\sigma_{\rm Tail}$ becomes non-trivial only when a large number of delays $(d=25)$ are taken and the model order is at-least 21. Consequently, Fig. 2 is a bona fide plot of Relative distance to DFT  for $d=3~{}\textrm{and}~{}12.$ Furthermore, the blue trend corresponding to $d=25$ depicts Relative distance to DFT  for $\theta\leq 20$, and an upper bound of the same for larger values of $\theta$.
Figure 5: Variation in $\sigma_{\rm Tail}$ for the computations with the Van der Pol oscillator in Fig. 2. Unless the model order is more than $20$ and $d=25$, $\sigma_{\rm Tail}$ remains low. Hence, the trends seen in Fig. 2, particularly the spike at $\theta=11$, are dynamically relevant and not an artifact of the low-rank approximation used in computing Relative distance to DFT.
Finally, Fig. 6 tells us that, with the exception of panel (d), Fig. 3 depicts only upper bounds on Relative distance to DFT. In panel (d) of Fig. 6, we see that $\sigma_{\rm Tail}$ is virtually zero for all choices of model order and delay embedding dimension. Hence, panel (d) of Fig. 3 is an authentic depiction of Relative distance to DFT. In contrast, panels (a), (b) and (c) see $\sigma_{\rm Tail}$ reach non-trivial values for both $d=12$ and $25$. Unfortunately, this departure from low discrepancy is the cause for panels (a) through (c) of Fig. 3 exhibiting a step-like trend. Despite this shortcoming, it is curious that said upper bounds are indicative of the growing complexity of the cavity flow.
(a) $Re~{}=~{}13\times 10^{3}$ (Periodic)
(b) $Re~{}=~{}16\times 10^{3}$ (Quasi-periodic)
(c) $Re~{}=~{}20\times 10^{3}$ (Mixed)
(d) $Re~{}=~{}30\times 10^{3}$ (Chaotic)
Figure 6: Variation of $\sigma_{\rm Tail}$ with model-order ($\theta$) and delay embedding dimension ($d$) for the cavity flow studied in Fig. 3. For Reynolds numbers between $13\times 10^{3}$ and $20\times 10^{3}$, $\sigma_{\rm Tail}$ becomes non-trivial even if only $12$ time delays are taken. Although $\sigma_{\rm Tail}$ has a far lower plateau than seen in the preceding dynamical systems, it is safer to interpret the concomitant computations in panels (a), (b) and (c) of Fig. 3 as upper bounds on Relative distance to DFT. The case of $Re~{}=~{}30\times 10^{3}$ is in stark contrast as $\sigma_{\rm Tail}$ is virtually non-existent for all choices of delay-embedding dimension ($d$) and model-order ($\theta$). Hence, panel (d) in Fig. 3 is a faithful representation of Relative distance to DFT.